OpenStudy (anonymous):

Please help What is the distance between the points (–3, 4) and (4, 6)? A.10 B.11 C.root of 53 D.root of 101

OpenStudy (anonymous):

C. root of 53 will be your answer

OpenStudy (paxpolaris):

\[\sqrt{\left( x_1-x_2 \right)^2+\left( y_1-y_2 \right)^2}\]

OpenStudy (anonymous):

What is the midpoint of a segment whose endpoints are (9, −9) and (−3, 7)? A. (3, −1) B. (6, 8) C. (−1, 3) D. (−8, 6)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

What is the midpoint of a segment whose endpoints are (9, −9) and (−3, 7)?

OpenStudy (anonymous):

sorry i reposted it

OpenStudy (paxpolaris):

take the average of the x coordinates and the average of the y coordinates \[\left( {9-3\over2},{-9+7\over2} \right)\] A